# Sequences with identical autocorrelation functions

### v2

@@ 1. Introduction
In many physical measurements of wave phenomena, detectors are unable to discern phases. This loss of phase information is called the phase problem, a terminology that arose in x-ray crystallography, where a diffraction pattern gives the magnitude of the Fourier transform of the electron density without the phase information [BE22]. Knowing the magnitude of the Fourier transform is the same as knowing the autocorrelation of the electron density, which in general is not sufficient information to recover the electron density itself. This paper concerns itself with the aperiodic one-dimensional discrete problem of phases, that is, the extent to which one can deduce a sequence from its aperiodic autocorrelation. See [BE22] on the the practical importance of the aperiodic problem of phases. Autocorrelation of sequences is important in many applications in communications and remote sensing where accurate timing and synchronization are required; see [Gol67, GG05].
@@
Because we consider aperiodic autocorrelation, a sequence is any doubly infinite sequence $f=(\ldots,f_{-1},f_{0},f_{1},f_{2},\ldots)$ of complex numbers such that only finitely many of the terms are nonzero. We identify this sequence with $f(z)=\sum_{j\in{\mathbb{Z}}}f_{j}z^{j}\in{\mathbb{C}}[z,z^{-1}]$, where ${\mathbb{C}}[z,z^{-1}]$ is the ring of Laurent polynomials with complex coefficients. If $R$ is any ring, then $R^{\times}$ denotes its group of units, and we say that two elements $f$ and $g$ are $R$-associates to mean that there is some $u\in R^{\times}$ such that $f=ug$. Notice that the units of ${\mathbb{C}}[z,z^{-1}]$ are monomials with nonzero coefficients, that is, elements of the form $cz^{j}$ with $j\in{\mathbb{Z}}$ and $c\in{\mathbb{C}}^{\times}$. Multiplication by a unit in the Laurent polynomial formalism shifts and scales a sequence, which for our purposes produces an equivalent sequence, so we are only interested in sequences up to the relation of being ${\mathbb{C}}[z,z^{-1}]$-associates.
The support of a sequence $f$, written $\operatorname{supp}f$, is the set $\{j\in{\mathbb{Z}}:f_{j}\not=0\}$. A segment is a set of consecutive integers. The length of a sequence $f$, written $\operatorname{len}f$, is the cardinality of the smallest segment that contains $\operatorname{supp}f$. A contiguous sequence $f$ is a sequence where $\operatorname{supp}f$ is a segment. For any positive integer $m$, an $m$-ary sequence is a contiguous sequence where $f_{j}$ is an $m$th root of unity in ${\mathbb{C}}$ for every $j\in\operatorname{supp}f$; when $m=2$, we have a binary sequence, where $f_{j}\in\{1,-1\}$ for every $j\in\operatorname{supp}f$.
@@
Note that the finite support of $f$ guarantees that $C_{f}(s)\not=0$ for only finitely many $s\in{\mathbb{Z}}$. The Laurent polynomial interpretation of sequences provides a convenient formalism for calculating autocorrelation by thinking of sequences as functions on the complex unit circle. To use this formalism, for $f(z)=\sum_{j\in{\mathbb{Z}}}f_{j}z^{j}\in{\mathbb{C}}[z,z^{-1}]$ we define the conjugate of $f(z)$, written $\overline{f(z)}$, to be $\sum_{j\in{\mathbb{Z}}}\overline{f_{j}}z^{-j}$. (Notice that conjugation of a sequence reverses the order of its elements and then replaces each one with its complex conjugate.) A self-conjugate element $f(z)$ of ${\mathbb{C}}[z,z^{-1}]$ is one for which $\overline{f(z)}=f(z)$. With conjugation defined, one readily shows that
@@
For $f\in{\mathbb{C}}[z,z^{-1}]$, the associate class of $f$, written $[{f}]$, is the set of all ${\mathbb{C}}[z,z^{-1}]$-associates of $f$, so $[{f}]=\{cz^{j}f(z):c\in{\mathbb{C}}^{\times},j\in{\mathbb{Z}}\}$. The trivial equicorrelationality class of $f$, written $[\kern-1.5pt[{f}]\kern-1.5pt]$, is the set of all sequences that are trivially equicorrelational to $f$, so $[\kern-1.5pt[{f}]\kern-1.5pt]=[f]\cup[\overline{f}]$. The equicorrelationality class of $f$, written $[\kern-1.5pt[\kern-1.5pt[{f}]\kern-1.5pt]\kern-1.5pt]$, is the set of all sequences that are equicorrelational to $f$, and is a union of trivial equicorrelationality classes. If $F$ is a self-conjugate subfield of ${\mathbb{C}}$ and $f(z)\in F[z,z^{-1}]$, then the $F[z,z^{-1}]$-associate class of $f$, written $[{f}]_{F}$, is the set of all associates of $f$ in $F[z,z^{-1}]$, which is just $[{f}]\cap F[z,z^{-1}]$ (see Lemma 2.5 for a proof). The $F$-trivial equicorrelationality class of $f$, written $[\kern-1.5pt[{f}]\kern-1.5pt]_{F}$, is the set of all sequences in $F[z,z^{-1}]$ that are trivially equicorrelational to $f$, that is, $[\kern-1.5pt[{f}]\kern-1.5pt]\cap F[z,z^{-1}]$, which is equal to $[{f}]_{F}\cup[{\overline{f}}]_{F}$ (see Lemma 2.12 for a proof). The $F$-equicorrelationality class of $f$, written $[\kern-1.5pt[\kern-1.5pt[{f}]\kern-1.5pt]\kern-1.5pt]_{F}$, is the set of all sequences in $F[z,z^{-1}]$ that are equicorrelational to $f$, that is, $[\kern-1.5pt[\kern-1.5pt[{f}]\kern-1.5pt]\kern-1.5pt]\cap F[z,z^{-1}]$.
A generalized palindrome is a sequence $f(z)$ in ${\mathbb{C}}[z,z^{-1}]$ that is a ${\mathbb{C}}[z,z^{-1}]$-associate of its own conjugate, that is, $\overline{f}\in[{f}]$. If $F$ is a self-conjugate subfield of ${\mathbb{C}}$, each $F[z,z^{-1}]$-associate class either consists entirely of generalized palindromes and is self-conjugate, or else the class has no generalized palindromes and its conjugate is a different $F[z,z^{-1}]$-associate class (see Lemma 2.11 for a proof).
@@
(1) $f(z)=uf_{1}^{a_{1}}\cdots f_{m}^{a_{m}}g_{1}^{b_{1}}\cdots g_{n}^{b_{n}}\overline{g_{1}}^{c_{1}}\cdots\overline{g_{n}}^{c_{n}}$
@@
Furthermore, we show that certain kinds of generalized palindromes cannot be equicorrelational to each other. A palindrome is a sequence $f\in{\mathbb{R}}[z,z^{-1}]$ such that $\overline{f(z)}=z^{j}f(z)$ for some $j\in{\mathbb{Z}}$. An antipalindrome is a sequence $f\in{\mathbb{R}}[z,z^{-1}]$ such that $\overline{f(z)}=-z^{j}f(z)$ for some $j\in{\mathbb{Z}}$. Palindromes and antipalindromes are the only kinds of generalized palindromes that occur among the binary sequences, and the only sequence that is both a palindrome and an antipalindrome is the zero sequence. We shall prove the following as a consequence of Theorem 1.1.
@@
For the rest of this introduction, we restrict the relation of equicorrelationality to binary sequences: an equivalence class of this relation is called a binary equicorrelationality class. We also restrict the notion of trivial equicorrelationality to binary sequences: two binary sequences $f$ and $g$ are trivially equicorrelational if and only if $f(z)=uz^{j}g(z)$ or $f(z)=uz^{j}\overline{g(z)}$ for some $u\in\{-1,1\}$ and $j\in{\mathbb{Z}}$, and an equivalence class of this relation is called a trivial binary equicorrelationality class. Trivial binary equicorrelationality refines binary equicorrelationality, so every binary equicorrelationality class is a union of pairwise disjoint trivial binary equicorrelationality classes. The volume of a binary equicorrelationality class equals the number of trivial binary equicorrelationality classes in this union, and a binary equicorrelationality class with volume greater than one is called nontrivial.
@@ 2. Preliminaries
Throughout this paper, =${\mathbb{N}}=\{0,1,2,\ldots\}$ and if $R$ is a ring, then $R^{\times}$ denotes the unit group of $R$. We always use the Laurent polynomial formalism for sequences and their autocorrelation functions, as described in the Introduction, so sequences are always thought of as elements of the Laurent polynomial ring ${\mathbb{C}}[z,z^{-1}]$. For any field $F$, we have $F[z,z^{-1}]^{\times}=\{cz^{j}:c\in F^{\times},j\in{\mathbb{Z}}\}$, so every nonzero $f\in F[z,z^{-1}]$ can be written uniquely as $f=ug$ where $u\in F[z,z^{-1}]^{\times}$ and $g$ is a monic polynomial in $F[z]$ with a nonzero constant coefficient, and then the length (originally described in the third paragraph of the Introduction) of $f$ is $1+\deg g$. (And, of course, $\operatorname{len}(0)=0$.) In particular, an element of $F[z,z^{-1}]$ is a unit if and only if it has length $1$. Then one can verify that $F[z,z^{-1}]$ is a Euclidean domain with $\operatorname{len}$ as its Euclidean size function. In particular, $\operatorname{len}(fg)\geq\operatorname{len}f$ for all $f,g\in F[z,z^{-1}]$ with $g\not=0$, and thus $F[z,z^{-1}]$-associates have the same length. Since $F[z,z^{-1}]$ is a Euclidean domain, every pair of elements $f,g\in F[z,z^{-1}]$ has a greatest common divisor (defined up to $F[z,z^{-1}]$-associates), and if $f$ and $g$ are not both 00, then we shall use $\gcd(f,g)$ to denote the unique monic polynomial with nonzero constant coefficient in the $F[z,z^{-1}]$-associate class containing the greatest common divisors (and we set $\gcd(0,0)=0$).
@@
In the Introduction, it was claimed that if $F$ is a subfield of ${\mathbb{C}}$ and $f(z)\in F[z,z^{-1}]$, then $[{f}]_{F}=[{f}]\cap F[z,z^{-1}]$, that is, the set of $F[z,z^{-1}]$-associates of $f(z)$ is obtained from the set $[{f}]$ of ${\mathbb{C}}[z,z^{-1}]$-associates of $f(z)$ by just taking those elements in the latter set whose coefficients all lie in $F$. We prove a slightly more general result here (where ${\mathbb{C}}$ is replaced with an arbitrary extension field $E$ of $F$).
@@
Now suppose that $f$ and $g$ are $E[z,z^{-1}]$-associates. If one of $f$ or $g$ is zero, then the other must also be zero, and then they are clearly $F[z,z^{-1}]$-associates. So we may assume that $f$ and $g$ are nonzero from now on, and then there is some unit $u$ of $E[z,z^{-1}]$ such that $f=ug$. Now $u(z)=ez^{j}$ for some nonzero $e\in E$. If $f_{k}z^{k}$ is the lowest degree monomial in $f$ (so $f_{k}\in F^{\times}$), then $ef_{k}z^{k+j}$ is the lowest degree monomial in $g$ (so $ef_{k}\in F^{\times}$). Thus $e=(ef_{k})/f_{k}\in F^{\times}$. This makes $u$ a unit in $F[z,z^{-1}]$, and so $f$ and $g$ are $F[z,z^{-1}]$-associates. This proves the first claim in this lemma.
@@
Let $f$ and $g$ be generalized palindromes with $\overline{f(z)}=uz^{j}f(z)$ and $\overline{g(z)}=vz^{k}g(z)$ for some $u,v\in{\mathbb{C}}^{\times}$ and $j,k\in{\mathbb{Z}}$. Let $F$ be a subfield of ${\mathbb{C}}$ and suppose that $f,g\in F[z,z^{-1}]$ and that $f$ and $g$ are ${\mathbb{C}}[z,z^{-1}]$-associates. Then either (i) $f=g=0$ or else (ii) $v/u=\overline{w}/w$ for some $w\in F^{\times}$ and $j\equiv k\pmod{2}$.
@@
Since $f$ and $g$ are ${\mathbb{C}}[z,z^{-1}]$-associates, either $f=g=0$ or both $f$ and $g$ are nonzero; we are done in the former case, so assume that the latter case holds. Then there is some $w\in{\mathbb{C}}^{\times}$ and $i\in{\mathbb{Z}}$ such that $g(z)=wz^{i}f(z)$. We conjugate both sides to obtain $\overline{g(z)}=\overline{w}z^{-i}\overline{f(z)}$ and substitute the expressions for $\overline{g(z)}$ and $\overline{f(z)}$ from the statement of the lemma to obtain $vz^{k}g(z)=\overline{w}z^{-i}uz^{j}f(z)$, and then replace $g(z)$ with $wz^{i}f(z)$ again to obtain $vz^{k}wz^{i}f(z)=\overline{w}z^{-i}uz^{j}f(z)$. Since $F[z,z^{-1}]$ is an integral domain, we can cancel out the nonzero term $f(z)$ to obtain $vwz^{i+k}=u\overline{w}z^{j-i}$, and matching constants and exponents produces $v/u=\overline{w}/w$ and $j=k+2i$. ∎

### v3

@@ 1. Introduction
In many physical measurements of wave phenomena, detectors are unable to discern phases. This loss of phase information is called the phase problem, a terminology that arose in x-ray crystallography, where a diffraction pattern gives the magnitude of the Fourier transform of the electron density without the phase information [BE22]. Knowing the magnitude of the Fourier transform is the same as knowing the autocorrelation of the electron density, which in general is not sufficient information to recover the electron density itself. Classical x-ray crystallography provides the periodic autocorrelation of electron density of the contents of a unit cell of a crystal. Modern imaging techniques have prompted researchers to also investigate phase retrieval in the aperiodic regime; see [BE22, pp. 1491–1492] and [SEC+15]. This paper concerns itself with the aperiodic one-dimensional discrete problem of phases, that is, the extent to which one can deduce a sequence from its aperiodic autocorrelation. Autocorrelation of sequences is important in many applications in communications and remote sensing where accurate timing and synchronization are required; see [Gol67, GG05]. As some examples, many foundational digital communications protocols such as code-division multiple access (CDMA) and orthogonal frequency-division multiplexing (OFDM) use low autocorrelation sequences, as do pulse compression schemes for efficient operation of radar. When a CDMA system uses a sequence for modulation, the sequence’s aperiodic autocorrelation determines its periodic and negaperiodic (also known as odd periodic) correlation functions, both which are important in determining the performance of the system [SP80, Sec. V.B]. Therefore, aperiodic autocorrelation can be viewed as the central object of interest in such systems.
@@
Because we consider aperiodic autocorrelation, a sequence is any doubly infinite sequence $f=(\ldots,f_{-1},f_{0},f_{1},f_{2},\ldots)$ of complex numbers such that only finitely many of the terms are nonzero. We identify this sequence with $f(z)=\sum_{j\in{\mathbb{Z}}}f_{j}z^{j}\in{\mathbb{C}}[z,z^{-1}]$, where ${\mathbb{C}}[z,z^{-1}]$ is the ring of Laurent polynomials with complex coefficients. Whenever we simply write a letter like “$g$” for a Laurent polynomial, it should be interpreted as shorthand for “$g(z)$”. Sometimes we write the full “$g(z)$” notation, especially when distinguishing $g(z)$ from other polynomials derived from $g(z)$ such as $g(-z)$ or $g(z^{2})$. If $R$ is any ring, then $R^{\times}$ denotes its group of units, and we say that two elements $f$ and $g$ are $R$-associates to mean that there is some $u\in R^{\times}$ such that $f=ug$. Notice that the units of ${\mathbb{C}}[z,z^{-1}]$ are monomials with nonzero coefficients, that is, elements of the form $cz^{j}$ with $j\in{\mathbb{Z}}$ and $c\in{\mathbb{C}}^{\times}$. Multiplication by a unit in the Laurent polynomial formalism shifts and scales a sequence, which for our purposes produces an equivalent sequence, so we are only interested in sequences up to the relation of being ${\mathbb{C}}[z,z^{-1}]$-associates.
The support of a sequence $f$, written $\operatorname{supp}f$, is the set $\{j\in{\mathbb{Z}}:f_{j}\not=0\}$. A segment is a set of consecutive integers. The length of a sequence $f$, written $\operatorname{len}f$, is the cardinality of the smallest segment that contains $\operatorname{supp}f$. A contiguous sequence $f$ is a sequence where $\operatorname{supp}f$ is a segment. For any positive integer $m$, an $m$-ary sequence is a contiguous sequence where $f_{j}$ is an $m$th root of unity in ${\mathbb{C}}$ for every $j\in\operatorname{supp}f$; when $m=2$, we have a binary sequence, where $f_{j}\in\{1,-1\}$ for every $j\in\operatorname{supp}f$. For readers more familiar with considering a binary sequence as a vector of 00s and $1$s from the binary finite field ${\mathbb{Z}}/2{\mathbb{Z}}$, we remark that we can represent such a sequence as a vector of $+1$s and $-1$s in ${\mathbb{C}}$ by applying the transformation $x\mapsto(-1)^{x}$ to the terms. Correlation measures the resemblance between two binary vectors by counting the number of coordinates where the vectors agree and deducting the number of coordinates where they disagree, so using the +1/1+1/-1+ 1 / - 1 representation of binary sequences makes correlation equal to the dot product of the two vectors. See [Gol94, p. 14] for how the discussion of binary sequences passes naturally from $0/1$ representation to +1/1+1/-1+ 1 / - 1 representation, and [SP80, pp. 595–596] on why more general sequences whose terms lie in ${\mathbb{C}}$ are considered in communications systems. For these more general sequences, correlation is still a dot product calculated in ${\mathbb{C}}$. It is important that the correlation calculation happens in a ring of characteristic 00 (such as ${\mathbb{C}}$) and not in a finite ring (such as ${\mathbb{Z}}/2{\mathbb{Z}}$, in which the terms of $0/1$-binary sequences lie) because finite rings have modular arithmetic, which would result in correlation values vanishing due to modular reduction in many cases where there is substantial agreement between sequences. For example, $m$-ary sequences can be considered as vectors of elements of ${\mathbb{Z}}/m{\mathbb{Z}}$, but for the purposes of calculating correlation, one uses the map $x\mapsto\exp(2\pi ix/m)$ to transform them into sequences whose terms are complex $m$th roots of unity.
@@
Note that the finite support of $f$ guarantees that $C_{f}(s)\not=0$ for only finitely many $s\in{\mathbb{Z}}$. The Laurent polynomial interpretation of sequences provides a convenient formalism for calculating autocorrelation by thinking of sequences as functions on the complex unit circle. To use this formalism, for $f(z)=\sum_{j\in{\mathbb{Z}}}f_{j}z^{j}\in{\mathbb{C}}[z,z^{-1}]$ we define the conjugate of $f(z)$, written $\overline{f(z)}$, to be $\sum_{j\in{\mathbb{Z}}}\overline{f_{j}}z^{-j}$. (Notice that conjugation of a sequence reverses the order of its terms and then replaces each one with its complex conjugate.) A self-conjugate element $f$ of ${\mathbb{C}}[z,z^{-1}]$ is one for which $\overline{f}=f$. With conjugation defined, one readily shows that
@@
For $f\in{\mathbb{C}}[z,z^{-1}]$, the associate class of $f$, written $[{f}]$, is the set of all ${\mathbb{C}}[z,z^{-1}]$-associates of $f$, so $[{f}]=\{cz^{j}f:c\in{\mathbb{C}}^{\times},j\in{\mathbb{Z}}\}$. The trivial equicorrelationality class of $f$, written $[\kern-1.5pt[{f}]\kern-1.5pt]$, is the set of all sequences that are trivially equicorrelational to $f$, so $[\kern-1.5pt[{f}]\kern-1.5pt]=[f]\cup[\overline{f}]$. The equicorrelationality class of $f$, written $[\kern-1.5pt[\kern-1.5pt[{f}]\kern-1.5pt]\kern-1.5pt]$, is the set of all sequences that are equicorrelational to $f$, and is a union of trivial equicorrelationality classes. If $F$ is a self-conjugate subfield of ${\mathbb{C}}$ and $f\in F[z,z^{-1}]$, then the $F[z,z^{-1}]$-associate class of $f$, written $[{f}]_{F}$, is the set of all associates of $f$ in $F[z,z^{-1}]$, which is just $[{f}]\cap F[z,z^{-1}]$ (see Lemma 2.5 for a proof). The $F$-trivial equicorrelationality class of $f$, written $[\kern-1.5pt[{f}]\kern-1.5pt]_{F}$, is the set of all sequences in $F[z,z^{-1}]$ that are trivially equicorrelational to $f$, that is, $[\kern-1.5pt[{f}]\kern-1.5pt]\cap F[z,z^{-1}]$, which is equal to $[{f}]_{F}\cup[{\overline{f}}]_{F}$ (see Lemma 2.12 for a proof). The $F$-equicorrelationality class of $f$, written $[\kern-1.5pt[\kern-1.5pt[{f}]\kern-1.5pt]\kern-1.5pt]_{F}$, is the set of all sequences in $F[z,z^{-1}]$ that are equicorrelational to $f$, that is, $[\kern-1.5pt[\kern-1.5pt[{f}]\kern-1.5pt]\kern-1.5pt]\cap F[z,z^{-1}]$.
A generalized palindrome is a sequence $f$ in ${\mathbb{C}}[z,z^{-1}]$ that is a ${\mathbb{C}}[z,z^{-1}]$-associate of its own conjugate, that is, $\overline{f}\in[{f}]$. If $F$ is a self-conjugate subfield of ${\mathbb{C}}$, each $F[z,z^{-1}]$-associate class either consists entirely of generalized palindromes and is self-conjugate, or else the class has no generalized palindromes and its conjugate is a different $F[z,z^{-1}]$-associate class (see Lemma 2.11 for a proof).
@@
(1) $f=uf_{1}^{a_{1}}\cdots f_{m}^{a_{m}}g_{1}^{b_{1}}\cdots g_{n}^{b_{n}}\overline{g_{1}}^{c_{1}}\cdots\overline{g_{n}}^{c_{n}}$
@@
Furthermore, we show that certain kinds of generalized palindromes cannot be equicorrelational to each other. A palindrome is a sequence $f\in{\mathbb{R}}[z,z^{-1}]$ such that $\overline{f}=z^{j}f$ for some $j\in{\mathbb{Z}}$. An antipalindrome is a sequence $f\in{\mathbb{R}}[z,z^{-1}]$ such that $\overline{f}=-z^{j}f$ for some $j\in{\mathbb{Z}}$. Palindromes and antipalindromes are the only kinds of generalized palindromes that occur among the binary sequences, and the only sequence that is both a palindrome and an antipalindrome is the zero sequence. We shall prove the following as a consequence of Theorem 1.1.
@@
For the rest of this introduction, we restrict the relation of equicorrelationality to binary sequences: an equivalence class of this relation is called a binary equicorrelationality class. We also restrict the notion of trivial equicorrelationality to binary sequences: two binary sequences $f$ and $g$ are trivially equicorrelational if and only if $f=uz^{j}g$ or $f=uz^{j}\overline{g}$ for some $u\in\{-1,1\}$ and $j\in{\mathbb{Z}}$, and an equivalence class of this relation is called a trivial binary equicorrelationality class. Trivial binary equicorrelationality refines binary equicorrelationality, so every binary equicorrelationality class is a union of pairwise disjoint trivial binary equicorrelationality classes. The volume of a binary equicorrelationality class equals the number of trivial binary equicorrelationality classes in this union, and a binary equicorrelationality class with volume greater than one is called nontrivial.
@@ 2. Preliminaries
Throughout this paper, =${\mathbb{N}}=\{0,1,2,\ldots\}$ and if $R$ is a ring, then $R^{\times}$ denotes the unit group of $R$. We always use the Laurent polynomial formalism for sequences and their autocorrelation functions, as described in the Introduction, so sequences are always thought of as elements of the Laurent polynomial ring ${\mathbb{C}}[z,z^{-1}]$. We retain the convention that whenever we simply write a letter like “$g$” for a Laurent polynomial, it should be interpreted as shorthand for “$g(z)$”, but we sometimes write the full “$g(z)$” notation, especially when distinguishing $g(z)$ from other polynomials derived from $g(z)$ such as $g(-z)$ or $g(z^{2})$. For any field $F$, we have $F[z,z^{-1}]^{\times}=\{cz^{j}:c\in F^{\times},j\in{\mathbb{Z}}\}$, so every nonzero $f\in F[z,z^{-1}]$ can be written uniquely as $f=ug$ where $u\in F[z,z^{-1}]^{\times}$ and $g$ is a monic polynomial in $F[z]$ with a nonzero constant coefficient, and then the length (originally described in the third paragraph of the Introduction) of $f$ is $1+\deg g$. (And, of course, $\operatorname{len}(0)=0$.) In particular, an element of $F[z,z^{-1}]$ is a unit if and only if it has length $1$. Then one can verify that $F[z,z^{-1}]$ is a Euclidean domain with $\operatorname{len}$ as its Euclidean size function. In particular, $\operatorname{len}(fg)\geq\operatorname{len}f$ for all $f,g\in F[z,z^{-1}]$ with $g\not=0$, and thus $F[z,z^{-1}]$-associates have the same length. Since $F[z,z^{-1}]$ is a Euclidean domain, every pair of elements $f,g\in F[z,z^{-1}]$ has a greatest common divisor (defined up to $F[z,z^{-1}]$-associates), and if $f$ and $g$ are not both 00, then we shall use $\gcd(f,g)$ to denote the unique monic polynomial with nonzero constant coefficient in the $F[z,z^{-1}]$-associate class containing the greatest common divisors (and we set $\gcd(0,0)=0$).
@@
In the Introduction, it was claimed that if $F$ is a subfield of ${\mathbb{C}}$ and $f\in F[z,z^{-1}]$, then $[{f}]_{F}=[{f}]\cap F[z,z^{-1}]$, that is, the set of $F[z,z^{-1}]$-associates of $f$ is obtained from the set $[{f}]$ of ${\mathbb{C}}[z,z^{-1}]$-associates of $f$ by just taking those elements in the latter set whose coefficients all lie in $F$. We prove a slightly more general result here (where ${\mathbb{C}}$ is replaced with an arbitrary extension field $E$ of $F$).
@@
Now suppose that $f$ and $g$ are $E[z,z^{-1}]$-associates. If one of $f$ or $g$ is zero, then the other must also be zero, and then they are clearly $F[z,z^{-1}]$-associates. So we may assume that $f$ and $g$ are nonzero from now on, and then there is some unit $u$ of $E[z,z^{-1}]$ such that $f=ug$. Now $u=ez^{j}$ for some nonzero $e\in E$. If $f_{k}z^{k}$ is the lowest degree monomial in $f$ (so $f_{k}\in F^{\times}$), then $ef_{k}z^{k+j}$ is the lowest degree monomial in $g$ (so $ef_{k}\in F^{\times}$). Thus $e=(ef_{k})/f_{k}\in F^{\times}$. This makes $u$ a unit in $F[z,z^{-1}]$, and so $f$ and $g$ are $F[z,z^{-1}]$-associates. This proves the first claim in this lemma.
@@
Let $f$ and $g$ be generalized palindromes with $\overline{f}=uz^{j}f$ and $\overline{g}=vz^{k}g$ for some $u,v\in{\mathbb{C}}^{\times}$ and $j,k\in{\mathbb{Z}}$. Let $F$ be a subfield of ${\mathbb{C}}$ and suppose that $f,g\in F[z,z^{-1}]$ and that $f$ and $g$ are ${\mathbb{C}}[z,z^{-1}]$-associates. Then either (i) $f=g=0$ or else (ii) $v/u=\overline{w}/w$ for some $w\in F^{\times}$ and $j\equiv k\pmod{2}$.
@@
Since $f$ and $g$ are ${\mathbb{C}}[z,z^{-1}]$-associates, either $f=g=0$ or both $f$ and $g$ are nonzero; we are done in the former case, so assume that the latter case holds. Then there is some $w\in{\mathbb{C}}^{\times}$ and $i\in{\mathbb{Z}}$ such that $g=wz^{i}f$. We conjugate both sides to obtain $\overline{g}=\overline{w}z^{-i}\overline{f}$ and substitute the expressions for $\overline{g}$ and $\overline{f}$ from the statement of the lemma to obtain $vz^{k}g=\overline{w}z^{-i}uz^{j}f$, and then replace $g$ with $wz^{i}f$ again to obtain $vz^{k}wz^{i}f=\overline{w}z^{-i}uz^{j}f$. Since $F[z,z^{-1}]$ is an integral domain, we can cancel out the nonzero term $f$ to obtain $vwz^{i+k}=u\overline{w}z^{j-i}$, and matching constants and exponents produces $v/u=\overline{w}/w$ and $j=k+2i$. ∎